Suppose National Collegiate Athletic Association [NCAA] rules state all student-athletes are to receive an average of 50 hours of academic support, per term. A random sample of 49 University of Maryland student-athletes revealed a standard deviation of 10 hours of academic support per term. If the calculated value for the associated test statistic equaled -1.75, what was the mean number of hours of academic support the student-athletes in the sample received per term? Group of answer choices 67.5 52.5 47.5 32.5
Answer #1

Solution :

= t * s+ = -1.75 * 10/sqrt49 + 50

= -1.75 * (10/7) + 50

= 47.5

Mean = 47.5